please help me answer this question. its from "solving systems by substitution" they just didn't have this option.
3x+2y=2
y=x-4

Answer by stanbon(75887)

3x+2y=2
y=x-4
-------
Substitute for "y" and solve for "x":
3x + 2(x-4) = 2
3x + 2x - 8 = 2
----
5x = 10
x = 2
-------------
Cheers,
Stan H.
=============